1.) At the end of the most recent year, Red's Cleaning Service had net income of $2,310. Red's Cleaning Service cost of goods sold was $3,850 and depreciation expense was $1,210. In addition, the firm paid $1,060 in interest expense, $1,110 in taxes, and $450 in dividends. Calculate Red's Cleaning Service sales. DO NOT USE DOLLAR SIGNS OR COMMAS IN YOUR ANSWER. ROUND TO THE NEAREST DOLLAR (e.g., 8000 NOT $8,000).

2.) A firm has net working capital of $700. Long-term debt is $4,300, total assets are $7,600, and fixed assets are $4,100. What is the amount of the total liabilities?
Answer #1

1) Net Income = Sales - cost of goods sold - depreciation - interest expense - taxes

2,310 = Sales - 3,850 - 1,210 - 1,060 - 1,110

Sales = 2,310 + 7230

Sales = 9540

Note that dividend is not part of Net Income\

2) Current assets = total assets - fixed assets

Current assets = 7,600 - 4,100 = $3,500

Net working capital = current assets - current liabilities

700 = 3,500 - current liabilities

current liabilities = 3500 - 700

current liabilities = 2,800

Total liabilities = current liabilities + long-term debt

Total liabilities = 2,800 + 4,300

Total liabilities = $7,100

Can you please upvote? Thank You :-)


Can you please upvote? Thank You :-)